# One-Shot Method for Computing Generalized Winding Numbers

### v1

@@ Abstract
The generalized winding number is an essential part of the geometry processing toolkit, allowing to quantify how much a given point is inside a surface, often represented by a mesh or a point cloud, even when the surface is open, noisy, or non-manifold. Parameterized surfaces, which often contain intentional and unintentional gaps and imprecisions, would also benefit from a generalized winding number. Standard methods to compute it, however, rely on a surface integral, challenging to compute without surface discretization, leading to loss of precision characteristic of parametric surfaces.
We propose an alternative method to compute a generalized winding number, based only on the surface boundary and the intersections of a single ray with the surface. For parametric surfaces, we show that all the necessary operations can be done via a Sum-of-Squares (SOS) formulation, thus computing generalized winding numbers without surface discretization with machine precision. We show that by discretizing only the boundary of the surface, this becomes an efficient method.
We demonstrate an application of our method to the problem of computing a generalized winding number of a surface represented by a curve network, where each curve loop is surfaced via Laplace equation. We use the Boundary Element Method to express the solution as a parametric surface, allowing us to apply our method without meshing the surfaces. As a bonus, we also demonstrate that for meshes with many triangles and a simple boundary, our method is faster than the hierarchical evaluation of the generalized winding number while still being precise.
We validate our algorithms theoretically, numerically, and by demonstrating a gallery of results on a variety of parametric surfaces and meshes, as well uses in a variety of applications, including voxelizations and boolean operations.
Figure 1: We propose a novel one-shot method to compute generalized winding numbers that does not require discretization of the surface. Our method only uses the boundary of the surface and a single ray-surface intersection test. For parametric surfaces (a, Coons patches), we thus can compute generalized winding number with machine precision without meshing or sampling the surface. Similarly, we compute generalized winding numbers for a curve network surfaced with minimal surfaces to high precision (b). Finally, our method can be also applied to meshes (c), where for some meshes it can improve the performance over standard methods while staying exact. Piggy Bank mesh with minor modifications by belch.
@@ 1 Introduction
For meshes and point clouds, even representing open, noisy, and non-manifold geometry, the notion of generalized winding number captures how much a point is inside the surface Jacobson et al. (2013); Barill et al. (2018). Generalized winding numbers are used in a variety of applications, including tetrahedral meshing Hu et al. (2018), surface reconstruction Xu et al. (2023), and medical imaging Becciu et al. (2011), to name a few.
The standard methods of computing generalized winding numbers calculate the projected area of the surface onto a unit sphere. For parametric surfaces, which often contain gaps (Fig. 2a,b) and thus also may benefit from generalized winding number, computing such projected area requires surface discretization. This discretization may defeat the purpose of having a high-precision parametric surface, as it can lead to a significant loss of precision (Fig. 2c,d).
@@
We propose an alternative method to compute generalized winding number addressing this challenge. Our main insight is that for any surface, with or without boundary, computing a generalized winding number at a point requires the surface boundary only and intersections of the surface with a single ray. We furthermore show that in typical scenarios, when processing multiple query points, one can use less than one ray per query point on average, significantly reducing time complexity.
Our method is generic and can be applied to various geometry representations. For parametric surfaces, our method allows us to compute the winding number to machine precision, bypassing surface discretization. We show that the key components of our method, including the ray-surface intersection, can be done exactly via a convex Sum-of-Squares formulation Marschner et al. (2021). In practice, if both precision and efficiency are important, we can discretize only the boundary of the parametric shape, and still get a high-precision result while being memory efficient and fast.
We demonstrate an application of our method to piecewise minimal surfaces on 3D curve networks, collections of 3D curves forming one or multiple loops (Fig. 1), that are easy to create in virtual or augmented reality (VR/AR) contexts Google (2023), and often used for quick modeling, ideation, and visualization Gryaditskaya et al. (2020) (Fig. 16). We apply our method to compute a winding number of piecewise minimal surfaces by expressing them in a parametric form via Boundary Element Method (BEM). As before, our method allows us to avoid discretizing these surfaces.
For meshes with few boundary edges, in typical application scenarios such as voxelization, our method can be significantly faster than the hierarchical evaluation of the winding number Jacobson et al. (2013). While an approximation can be computed more efficiently Barill et al. (2018), our computation is exact.
@@
We showcase our method on a wide variety of standard tasks for generalized winding numbers for all these representations (Sec. 7), including parametric Coons patches and Bézier triangles. We extensively validate our method by both formally showing it is equal to the generalized winding number and validating its numerical precision and performance in numerous experiments, as compared to Jacobson et al. (2013); Barill et al. (2018) (Sec. 7).
@@
  • a new method of computing generalized winding number based on this observation, enabling exact computations for parametric surfaces, fast computations for some meshes, and computations for minimal surfaces requiring no meshing.
Figure 2: Parametric surfaces often contain gaps and overlaps, either intentional for tight alignment after manufacturing (a) or unintentional due to modelling defects. At the same time, precision is paramount for parametric surfaces, especially when parts are designed to touch (b). Discretizing such surfaces is often undesired, as it may introduce new artifacts (c). Our method can compute winding numbers precisely without surface discretization (d). Truck by Toru Odazawa.
@@ 2 Related Work
The classical concept of a winding number, originally defined for a closed curve Meister (1769) and easily generalizable to closed surfaces, is one of the topics in complex analysis Needham (1997). Jacobson et al. (2013) generalized winding numbers to polylines and meshes, including the ones with boundary. They also propose an efficient algorithm to compute the generalized winding number that empirically has a time complexity of $O(F^{0.55})$, where $F$ is the number of triangles. Later, winding numbers have been further generalized to, among other representations, point clouds and triangle soups Barill et al. (2018). The method of Barill et al. (2018) can also serve as a very efficient approximation for meshes, achieving $O(\log F)$. Under simple assumptions, on meshes our algorithm has a time complexity between those two, while still being exact (Sec. 7). More importantly, these methods target discretized geometry; applying them to parametric surfaces leads to loss of precision (Fig. 9). Our method can be applied to parametric surfaces directly yielding exact winding numbers. Finally, Feng et al. (2023b) extended the definition to curves on discrete surfaces.
@@
Our method computes a generalized winding number of a surface in any representation that supports ray intersection queries, such as parametric surfaces or meshes. A recent work by Spainhour et al. (2024) offers a practical method of finding a generalized winding number for a rational curve in 2D; their technique cannot be easily extended into 3D. In contrast, we propose a general method for any surfaces or curves supporting ray intersections, including parametric surfaces in 2D and 3D. We are not aware any previous work considering generalized winding number of parametric surfaces.
@@
A very related concept is a solid angle, which for a surface is defined as generalized winding number (without normalization) modulo $4\pi$, so it does not measure the number of turns the surface makes around a point, nor does it change under a flip of surface orientation Feng et al. (2023a). We are focusing on the full winding number that includes the multiplicity and depends on the orientation. The solid angle can be similarly defined and effectively computed for a space curve Knöppel (2018); Chern and Ishida (2024); Binysh and Alexander (2018). Note that in 3D we always compute a winding number of a surface, albeit sometimes not discretized.
Surface Reconstruction from 3D Curve Networks
Inspired by recent progress in 3D curve networks creation Bae et al. (2008); Schmidt et al. (2009); Yu et al. (2021), including industry standard applications in VR Adobe (2023); Google (2023); Gravity Sketch (2023), a series of works addressed the rather difficult problem of surface reconstruction from 3D curve networks. Clean curve networks with precise connectivity are often first broken into oriented closed loops and surfaced Abbasinejad (2011); Zhuang et al. (2013). Each loop can be surfaced via minimization of a variety of energies Welch and Witkin (1994); Bobenko and Schröder (2005); Moreton and Séquin (1992), including Dirichlet energy Schneider and Kobbelt (2001) or Laplacian energy Botsch and Kobbelt (2004). Those two energies are particularly interesting, because their first-order optimality condition is a linear PDE, Laplace and biharmonic equations Bloor and Wilson (1990, 1989). Our method is compatible with this last family of methods, where Boundary Element Methods (BEM) apply. Rough curve networks are often surfaced as a whole Huang et al. (2019); Lin et al. (2022). The desired surfaces can be, depending on the domain, piecewise-smooth Yu et al. (2022), curvature-aligned Bessmeltsev et al. (2012); Pan et al. (2015) or simply smooth Huang et al. (2019); Nealen et al. (2007). These reconstruction methods, producing a mesh or a parametric surface, are all compatible with our winding number computation method.
Ray Casting
@@
Our work is influenced by works on ray casting to voxelize shapes Nooruddin and Turk (2003); Houston et al. (2003). The surfaces they focus on have no true boundary, albeit might be composed of multiple components due to noise, so they perform a inside-outside test; ours may have a boundary, so the definition of insidedness is not binary and is better captured by the notion of generalized winding number.
@@ 3 Background
We start by defining a generalized winding number for a smooth [^1] Both the theory and the method we propose apply equally to piecewise smooth surfaces. We only use smoothness for the simplicity of exposition. surface $M$ with a boundary $\partial M$. For brevity, later in the text we write ‘winding number’ instead of ‘generalized winding number’.
For a given surface $M$, the winding number at a point $p\in\mathbb{R}^{3}~{}\setminus~{}M$ is defined as the integral of the differential solid angle $d\Omega(p)$ subtended at $p$ Barill et al. (2018):
[…]
An equivalent formulation of the winding number, which we use at a starting point for our method, is the number of signed intersections of rays with the surface over all directions Jacobson et al. (2013):
[…]
where $S^{2}_{p}$ is a unit sphere centered at $p$, and $\chi(q)$ is the number of signed intersections between ray $\overrightarrow{pq}$ and $M$. By ‘signed intersections’ we mean that an intersection is counted as $+1$ when the ray is intersecting from the back of the surface, i.e., the ray is aligned with the surface normal, and as $-1$ when it is intersecting from the front. So formally,
@@ 4 Algorithm
The input to our algorithm is a manifold surface $M$, with or without boundary $\partial M$. We first describe a generic algorithm, then clarify its implementation for parametric and discrete surfaces in Sec. 5.1.
@@ 4.1 Computing the Winding Number
Our goal is to compute the winding number at a point $p\in\mathbb{R}^{3}$, i.e., evaluate the integral in Eq. 1 using the boundary $\partial M$. Computing $\chi(q)$ for a single $q$ includes finding all intersections of a ray $\overrightarrow{pq}$ with a potentially non-discretized surface — a rather expensive operation (Sec. 6.2). Therefore, evaluating the integral naïvely is infeasible, as it requires many such ray-surface intersections. Instead, our main observation is the following lemma, a corollary of a more general and technical proposition we prove in the Supplementary materials:
@@
$w_{M}(p)=\frac{1}{{\color[rgb]{0.0,0.0,0.00}4}\pi}\sum_{i}\textrm{Area}_{S^{2}_{p}}(A_{i})\chi_{i}.$ (3)
The technical detail ‘almost everywhere’ in the formulation of the lemma refers to the points $q$ where the rays $\overrightarrow{pq}$ are tangent to the surface at a contact point (see inset). The set of such points, however, has measure zero, so they do not affect the integral (see Supplementary).
The observation captured in this lemma significantly simplifies our algorithm. Now we can shoot one ray per region $A_{i}$, compute the number of signed intersections $\chi_{i}$, and sum their contributions multiplied by the region area $\textrm{Area}_{S^{2}_{p}}(A_{i})$. Considering that number of regions is typically small, this leads to an efficient algorithm: project the boundary onto the sphere (Fig. 4c), decompose the sphere surface into regions, compute their areas, shoot a single ray through each one and compute $\chi_{i}$ (Fig. 4d).
@@
If the curve network has no boundary, i.e., represents a closed manifold, then the sphere only has one region $A_{1}$, so we compute only one $\chi_{1}$. Such manifolds divide the space into ‘inside’ and ‘outside’ (see Jordan–Brouwer separation theorem Guillemin and Pollack (2010)), so in this case our algorithm is equivalent to a simple inside-outside ray casting test.
@@ 5 One-shot algorithm
Performing ray-surface intersection tests for a non-discretized surface is still rather expensive, so we would like to minimize their number. Our key observation enabling an efficient algorithm is the following lemma:
@@
Indeed, crossing the projection of the boundary $\partial M$ means one of the regions has an additional intersection with $M$ (Fig. 6). This lemma is also a direct corollary of the proposition in the Supplementary. By ‘adjacent along a common edge’ we mean that two regions touching only at a common point do not follow that property. For instance, in Fig. 6, while the $\chi=0$ and $1$ regions are adjacent across edges and thus follow the lemma, the $\chi=1$ and $\chi=-1$ regions only share a common point, so $|\chi_{i}-\chi_{j}|=2$.
@@
If we find at least one intersection, we additionally test whether the ray is tangent to the surface at one of those points. If it is, we ignore the ray, because it is a point of contact rather than an intersection, so the number of intersection points for this ray is unreliable, and choose a different random ray through the region. We test this by verifying $|\overrightarrow{pq}\cdot n|>\varepsilon$, where $n$ is the surface normal. For BEM, we use $\varepsilon=10^{-2}$; for parametric geometry, $\varepsilon=10^{-12}$.
Once all the pairwise differences $\chi_{ij}$ are known for all adjacent regions $i,j$, we can represent adjacency between the regions as a directed graph, where each region becomes a vertex, and each pair of adjacent regions $i,j$ gets two edges $i\rightarrow j$ and $j\rightarrow i$, associated with $\chi_{ij}$ and $\chi_{ji}$, respectively. Since it is a connected graph, knowing $\chi_{i}$ for a single ’seed’ vertex $i$ is enough to compute all the other $\chi_{j}$. We therefore use breadth-first search on the graph starting from $i$ and fill in all the other values of $\chi_{j}$, which completes our algorithm. Note that the typical number of regions, and therefore graph vertices, is fewer than 5, so this is essentially instantaneous.
Optimization
We further observe that once we know all the intersections along the ray $\overrightarrow{pq}$, we can immediately compute the winding numbers for all the points along that ray without any more intersection tests. More precisely, once we compute all the ray parameter values $t_{j}$ where it intersects the surface, we know $\chi$ for any point with parameter $t$ along that ray by only counting the intersections with $t_{j}>t$. Thus, for every point along that ray, we only need to perform the decomposition into spherical regions to compute its winding number.
@@
6:     $\{r_{k},t_{k},n_{k}\}\leftarrow\textrm{Intersect}(\overrightarrow{pq})$triangleright $r_{k}\in\Omega,t_{k}\geq 0,n_{k}\in\mathbb{R}^{3}$
7:     $\chi_{i}\leftarrow\sum_{k}\textrm{Sign}(n_{k}\cdot\overrightarrow{pq})$triangleright Eq. 2
@@
10:     return $\frac{1}{4\pi}\sum_{j}\chi_{j}\cdot Area(A_{j})$triangleright Eq. 3

### v2

@@ Abstract
The generalized winding number is an essential part of the geometry processing toolkit, allowing to quantify how much a given point is inside a surface, even when the surface has boundaries and noise. We propose a new universal method to compute a generalized winding number, based only on the surface boundary and the intersections of a single ray with the surface, supporting any oriented surface representations that support a ray intersection query. Due to the focus on the boundary, our algorithm has a unique set of properties. For 2D parametric curves, on a regular grid of query points, our method is up to $4\times$ faster than the current state of the art, maintaining the same precision. In 3D, our method can compute a winding number of a surface without discretizing it, including parametric surfaces. For some meshes with many triangles and a simple boundary, our method is faster than the hierarchical evaluation of the generalized winding number while still being precise. Similarly, on some parametric surfaces with a simple boundary, our method can be faster than adaptive quadrature. We validate our algorithms theoretically, numerically, and by demonstrating a gallery of results on a variety of parametric surfaces and meshes, as well uses in a variety of applications, including voxelizations and boolean operations. {CCSXML} <ccs2012> <concept> <concept_id>10010147.10010371.10010396.10010402</concept_id> <concept_desc>Computing methodologies Shape analysis</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010396.10010399</concept_id> <concept_desc>Computing methodologies Parametric curve and surface models</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
volume: 44issue: 5
@@ 1 Introduction
For meshes, parametric surfaces, and point clouds, even representing open, noisy, and non-manifold geometry, the notion of generalized winding number captures how much a point is inside the surface [JKS13, BDS18] (Fig. 1). Generalized winding numbers are used in a variety of applications, including tetrahedral meshing [HZG18], surface reconstruction [XDW23], and medical imaging [BFP11], to name a few.
@@
The standard methods of computing generalized winding numbers calculate the area of the surface projected onto a unit sphere. We propose an alternative method to compute the generalized winding number. Our main insight is that for any surface, with or without boundary, computing a generalized winding number at a point requires the surface boundary only and intersections of the surface with a single ray. We furthermore show that in typical scenarios, when processing multiple query points, one can use less than one ray per query point on average, significantly reducing time complexity.
Our method is generic and can be applied to various geometry representations, in 2D or 3D. The elegance lies in its universality: just like a standard ray tracing pipeline, our method is exactly the same for different surface representations, such as meshes or parametric surfaces, and encapsulates the representation completely within a ray-surface intersection test. Due to our method’s focus on the boundary, the performance of our method does not depend much on the complexity of the surface and instead excels when the boundary is simple. For example in 2D, when the boundary of a set of parametric curves is just a set of points, our method, when run on a regular grid of query points is up to 4x faster than the state of the art. Similarly, for challenging parametric surfaces with a planar boundary (Fig. 16), on a similar arrangement of query points, our method is significantly faster than the standard integration method, adaptive quadrature, while maintaining the same precision. For meshes with few boundary edges, in typical application scenarios such as voxelization, our method can be significantly faster than the hierarchical evaluation of the winding number [JKS13]. While an approximation can be computed more efficiently [BDS18], our computation is significantly more precise. The regular grid of query points allows us to reuse the same ray intersection for many points, saving computations. We additionally show that the key components of our method, including the ray-surface intersection, can be done exactly via a convex Sum-of-Squares formulation [MZPS21], making the computations exact if so desired.
@@
We showcase our method on a wide variety of standard tasks for generalized winding numbers for all these representations (Sec. 7), in 2D and 3D, including parametric Coons patches and Bézier curves and triangles. We extensively validate our method by both formally showing it is equal to the generalized winding number and validating its numerical precision and performance in numerous experiments, as compared to the state of the art methods including [JKS13, BDS18, SGW24] and adaptive quadrature (Sec. 7). An application of our method is to piecewise minimal surfaces on 3D curve networks, collections of 3D curves forming one or multiple loops (Fig. One-Shot Method for Computing Generalized Winding Numbers), typical for (VR/AR) contexts [Goo23, GHL20] (Fig. 10). We apply our method to compute a winding number of piecewise minimal surfaces by expressing them in a parametric form via Boundary Element Method (BEM). Our method allows us to avoid discretizing these surfaces.
@@
  • a new method of computing generalized winding number based on this observation, enabling state-of-the-art performance for 2D parametric geometry on a regular grid of query points, and showing performance improvements for some 3D parametric surfaces and meshes.
Figure 1: Just like meshes, parametric surfaces also often contain gaps and overlaps, either intentional for tight alignment after manufacturing (a) or unintentional due to modelling defects. At the same time, precision is paramount for parametric surfaces, especially when parts are designed to touch (b). Discretizing such surfaces is often undesired, as it may introduce new artifacts (c). Our method can compute winding numbers precisely without surface discretization (d). Truck by Toru Odazawa.
@@ 2 Related Work
Winding Numbers. The classical concept of a winding number, originally defined for a closed curve [Mei69] and easily generalizable to closed surfaces, is one of the topics in complex analysis [Nee97]. [JKS13] generalized winding numbers to polylines and meshes, including the ones with boundary. They also propose an efficient algorithm to compute the generalized winding number that empirically has a time complexity of $O(F^{0.55})$, where $F$ is the number of triangles. Later, winding numbers have been further generalized to, among other representations, point clouds and triangle soups [BDS18]. The method of [BDS18] can also serve as a very efficient approximation for meshes, achieving $O(\log F)$. Under simple assumptions, on meshes our algorithm has a time complexity between those two, while still being exact (Sec. 7). More importantly, these methods target discretized geometry; applying them to parametric surfaces leads to loss of precision (Fig. 15). Our method can be applied to parametric surfaces directly yielding exact winding numbers. Finally, [FGC23b] extended the definition to curves on discrete surfaces.
@@
Our method computes a generalized winding number of a surface in any representation that supports ray intersection queries, such as parametric surfaces or meshes. A recent work [SGW24] offers a practical method of finding a generalized winding number for a rational curve in 2D; their technique cannot be easily extended into 3D. In contrast, we propose a general method for any surfaces or curves supporting ray intersections, including parametric or discrete geometry in 2D and 3D. We compare with their method in 2D in Sec. 7.
@@
A closely related concept is a solid angle, which for a surface is defined as generalized winding number (without normalization) modulo $4\pi$, so it does not measure the number of turns the surface makes around a point, nor does it change under a flip of surface orientation [FGC23a]. We are focusing on the full winding number, the signed solid angle, that includes the multiplicity and depends on the orientation. The solid angle can be similarly defined and effectively computed for a space curve [Kno18, CI24, BA18]. Note that in 3D we always compute a winding number of a surface, albeit sometimes not discretized.
@@
Ray Casting. Our work is influenced by works on ray casting to voxelize shapes [NT03, HBW03]. The surfaces they focus on have no true boundary, albeit might be composed of multiple components due to noise, so they perform an inside-outside test; ours may have a boundary, so the definition of insidedness is not binary and is better captured by the notion of generalized winding number.
@@ 3 Background
We start by defining a generalized winding number for a smooth oriented curve $C$ in 2D and an oriented surface $M$ in 3D, open or closed. Both the theory and the method we propose apply equally to piecewise smooth curves or surfaces. We only use smoothness for the simplicity of exposition. For brevity, later in the text we write ‘winding number’ instead of ‘generalized winding number’.
Figure 2: In two dimensions, the winding number is the sum of the signed subtended angles of the curve.
In 2D, a winding number at a point $p\in\mathbb{R}^{2}\setminus C$ is defined via the subtended angle $\theta$ of the curve $C$:
$w_{C}(p)=\frac{1}{2\pi}\int_{C}{d\theta(p)}.$
In 3D, for a given surface $M$, the winding number at a point $p\in\mathbb{R}^{3}~\setminus~M$ is defined as the integral of the differential solid angle $d\Omega(p)$ subtended at $p$ [BDS18]:
$w_{M}(p)=\frac{1}{4\pi}\int_{M}{d\Omega(p)}.$
In the article we focus on the 3D case for brevity; our algorithm works in both 2D and 3D, as we demonstrate in Fig. One-Shot Method for Computing Generalized Winding Numbers and Sec. 7. In both the theory and the algorithm we consider the winding number to be undefined when $p\in C$ or $p\in M$.
An equivalent formulation of the winding number, which we use as the starting point for our method, is the number of signed intersections of rays with the surface over all directions [JKS13]:
$w_{M}(p)=\frac{1}{4\pi}\int_{q\in S^{2}_{p}}\chi(q)dA,$ (1)
where $S^{2}_{p}$ is a unit sphere centered at $p$, and $\chi(q)$ is the number of signed intersections between ray $\overrightarrow{pq}$ and $M$. By ‘signed intersections’ we mean that an intersection is counted as +1+1 when the ray is intersecting from the back of the surface, i.e., the ray is aligned with the surface normal, and as 1-1 when it is intersecting from the front. So formally,
@@ 4 Algorithm
The input to our algorithm is an oriented manifold surface $M$ of arbitrary genus (Fig. 3), with or without boundary $\partial M$. We first describe a generic algorithm, then clarify its implementation for parametric and discrete surfaces in Sec. 5.1.
Figure 3: Our method is compatible with oriented surfaces of arbitrary genus. We show a torus (a) with genus 1, and a (4,3)-torus knot with genus 3 (b). Surface boundaries are highlighted in black.
@@ 4.1 Computing the Winding Number
Our goal is to compute the winding number at a point $p\in\mathbb{R}^{3}$, i.e., evaluate the integral in Eq. 1 using the boundary $\partial M$. Computing $\chi(q)$ for a single $q$ includes finding all intersections of a ray $\overrightarrow{pq}$ with a potentially non-discretized surface — a rather expensive operation (Sec. 6.2). Therefore, evaluating the integral naïvely is infeasible, as it requires many such ray-surface intersections. Instead, our main observation is the following lemma, a corollary of a more general and technical proposition we prove in Appendix A:
@@
$w_{M}(p)=\frac{1}{4\pi}\sum_{i}\textrm{Area}_{S^{2}_{p}}(A_{i})\chi_{i}.$ (3)
The technical detail ’almost everywhere’ in the formulation of the lemma refers to the points $q$ where the rays $\overrightarrow{pq}$ are tangent to the surface at a contact point (see inset). The set of such points, however, has measure zero, so they do not affect the integral.
The observation captured in this lemma significantly simplifies our algorithm. Now we can shoot one ray per region $A_{i}$, compute the number of signed intersections $\chi_{i}$, and sum their contributions multiplied by the region area $\textrm{Area}_{S^{2}_{p}}(A_{i})$. Considering that number of regions is typically small, this leads to an efficient algorithm: project the boundary onto the sphere (Fig. 4c), decompose the sphere’s surface into regions, compute their areas, shoot a single ray through each one and compute $\chi_{i}$ (Fig. 4d). Therefore, our method is compatible with all oriented surface representations that support ray-intersection queries and represent boundaries as continuous curves. Examples of unsupported representations include oriented point clouds and NeRFs [MST21], as they neither provide explicit ray-intersections nor represent boundaries as continuous curves.
@@
If the curve network has no boundary, i.e., represents a closed manifold, then the sphere only has one region $A_{1}$, so we compute only one $\chi_{1}$. Such manifolds divide the space into ‘inside’ and ‘outside’ (see Jordan–Brouwer separation theorem [GP10]), so in this case our algorithm is equivalent to a simple inside-outside ray casting test.
@@ 5 One-shot algorithm
Performing ray-surface intersection tests is still rather expensive, so we would like to minimize their number. Our key observation enabling an efficient algorithm is the following lemma:
@@
Indeed, crossing the projection of the boundary $\partial M$ means one of the regions has an additional intersection with $M$ (Fig. 7). This lemma is also a direct corollary of the proposition in Appendix A. By ‘adjacent along a common edge’ we mean that two regions touching only at a common point do not follow that property. For instance, in Fig. 7, while the $\chi=0$ and $1$ regions are adjacent across edges and thus follow the lemma, the $\chi=1$ and $\chi=-1$ regions only share a common point, so $|\chi_{i}-\chi_{j}|=2$.
@@
If we find at least one intersection, we additionally test whether the ray is tangent to the surface at one of those points. If it is, we ignore the ray, because it is a point of contact rather than an intersection, so the number of intersection points for this ray is unreliable, and choose a different random ray through the region. We test this by verifying $|\overrightarrow{pq}\cdot n|>\varepsilon$, where $n$ is the surface normal. For BEM, we use $\varepsilon=10^{-2}$; for parametric geometry and meshes, $\varepsilon=10^{-12}$.
Once all the pairwise differences $\chi_{ij}$ are known for all adjacent regions $i,j$, with a similar approach to [ZGZJ16], we can represent adjacency between the regions as a directed graph, where each region becomes a vertex, and each pair of adjacent regions $i,j$ gets two edges $i\rightarrow j$ and $j\rightarrow i$, associated with $\chi_{ij}$ and $\chi_{ji}$, respectively. Since it is a connected graph, knowing $\chi_{i}$ for a single ’seed’ vertex $i$ is enough to compute all the other $\chi_{j}$. We therefore use breadth-first search on the graph starting from $i$ and fill in all the other values of $\chi_{j}$, which completes our algorithm. Note that the typical number of regions, and therefore graph vertices, is fewer than 5, so this is essentially instantaneous.
Optimization. We further observe that once we know all the intersections along the ray $\overrightarrow{pq}$, we can immediately compute the winding numbers for all the points along that ray without any more intersection tests. More precisely, once we compute all the ray parameter values $t_{j}$ where it intersects the surface, we know $\chi$ for any point with parameter $t$ along that ray by only counting the intersections with $t_{j}>t$. Thus, for every point along that ray, we only need to perform the decomposition into spherical regions to compute its winding number.
@@
6:  $\{r_{k},t_{k},n_{k}\}\leftarrow\textrm{Intersect}(\overrightarrow{pq})$ $\triangleright$ $r_{k}\in\Omega,t_{k}\geq 0,n_{k}\in\mathbb{R}^{3}$
7:  $\chi_{i}\leftarrow\sum_{k}\textrm{Sign}(n_{k}\cdot\overrightarrow{pq})$ $\triangleright$ Eq. 2
@@
10:  return $\frac{1}{4\pi}\sum_{j}\chi_{j}\cdot Area(A_{j})$ $\triangleright$ Eq. 3